OpenStudy (anonymous):

How many times does a lions heart beat in 60 seconds

OpenStudy (zzr0ck3r):

I'll be back with the answer in one minute.

OpenStudy (anonymous):

Depends on the activitiy of the lion... Is the lion active or fat and lazy

OpenStudy (anonymous):

45

OpenStudy (zzr0ck3r):

He tried to eat me, you are on your own for this one.

OpenStudy (anonymous):

lol ^

OpenStudy (anonymous):

if the lion is active the heart rate will be higher but if hes calm then itll be low

OpenStudy (anonymous):

u should of put the rest of the question

OpenStudy (anonymous):

A lion's heart beats 12 times in 16 seconds. How many heartbeats will it have in 60 seconds?

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

If thats the question then set \[\frac{ 12 }{ 16 }\frac{ x }{ 60 }\]it up as a proportion @zzr0ck3r doesnt look like a junkie lion \[\frac{ 12 }{ }\]

OpenStudy (anonymous):

ignore the 12 and division bar

OpenStudy (anonymous):

google it

OpenStudy (anonymous):

when yall is done can yall help me with a question i posted. thx

OpenStudy (anonymous):

What is the whole question?